Steam enters a counterflow heat exchanger operating at steady state at 0.07 MPa with a quality of 0.9 and exits at the same pressure as saturated liquid. The steam mass flow rate is 1.3 kg/min. A separate stream of air with a mass flow rate of 100 kg/min enters at 30oC and exits at 60oC. The ideal gas model with cp = 1.005 kJ/kg·K can be assumed for air. Kinetic and potential energy effects are negligible.

Determine the temperature of the entering steam, in oC.
For the overall heat exchanger as the control volume, what is the rate of heat transfer, in kW.
Answer #1

Steam enters counterflow heat exchanger at steady state P1 = 0.07 MPa and x = 0.9

Tsat = 89.93 C

saturated specific enthalpy of liquid hL = 376.68 kj/kg

Saturated specific enthalpy of vapor, hg = 2659.41 kj/kg

Enthalpy inlet ,

h1 = (1-x)*hL + xhg = (1-0.9)*376.68 + 0.9*2659.41

h1 = 2431.13 kj/kg.

Outlet steam condition : saturated liquid with same pressure P2= 0.07 MPa

Ts = T2 = 89.93 C

h2 = hL = 376.68 kj/kg

Heat transfer rate, : mass flow rate of steam m = 1.3kg/min = 1.3/60 kg/s = 0.02166 kg/s

Q = m(h1 - h2) =( 0.0216kg/s )*(2431.13 - 376.68) = 51kj/s = 51kW

a) heat transfer from steam to air = 51 kW

m(CpsteamT1 - CpliquidT2) = 51 kW

(0.0216kg/s)(2.041*T1 - 4.205*T2) = 51

2.041*T1 - 4.205*89.93 = 2361.11

T1 = 1342 C

Entering steam temperature T1 = 1342 C

B) rate of heat transfer Q = 51 kW